Find the value of tn–1,α needed to construct a 90% upper or lower confidence bound with sample size 12. Round the answer to three decimal places.
Answer #1

solution;

n = Degrees of freedom = df = n - 1 = 12 - 1 = 11

At 90% upper level the t is ,

= 1 - 90% = 1 - 0.90 = 0.1

t ,df = t0.1,11 = 1.363    ( using student t table)